OpenStudy (anonymous):

need to prove identities sin^2thetacos^2theta=1\8[1-cos(4theta)]

OpenStudy (anonymous):

LHS = (sin^2theta)(cos^2theta) = 0.5(1-cos2theta)(0.5)(cos2theta+1) = 0.25(1-cos^2 2theta) = 0.25(1-0.5(cos4theta+1)) = 1/8(1-cos4theta) = RHS please represent your answer in consistency when using fractions and decimals : )